OpenStudy (anonymous):

Simplify: 6xy^-1 _________ 2x^-2y^-1

OpenStudy (anonymous):

\[\frac{6xy^{-1}}{2x^{-2}y^{-1}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

divide 6 by 2 to get 3

OpenStudy (anonymous):

\[\frac{y^{-1}}{y^{-1}}=1\]

OpenStudy (anonymous):

okay 3 goes on top right?

OpenStudy (anonymous):

and finally \[\frac{x}{x^{-2}}=x\times x^2=x^3\]\]

OpenStudy (anonymous):

yes 3 goes on top

OpenStudy (anonymous):

so your "final answer" is \[3x^3\]

OpenStudy (anonymous):

thank you